OpenStudy (vera_ewing):

Is the answer 96 or 12?!

OpenStudy (vera_ewing):

What is the remainder when (2x3 + 4x2 - 32x + 18) ÷ (x + 3)?

OpenStudy (vera_ewing):

By the remainder theorem, when a polynomial f(x) is divided by (x - a), the remainder is f(a). Here we are dividing by (x - 3), so plug x = 3 into the numerator: r = f(3) r = 2(3)³ + 4(3)² - 32(3) + 18 r = 12

OpenStudy (vera_ewing):

But @amistre64 says the answer is 96?!

OpenStudy (anonymous):

just go with the answer you got first, then try his if yours is wrong or vise versa

OpenStudy (vera_ewing):

I need to know the right answer. Is it 12 or 96?

OpenStudy (anonymous):

i done the same thing you did and got the same answer,so go with 12

OpenStudy (vera_ewing):

Yeah but why would @amistre64 and @robtobey both say it's 96? Doesn't make sense to me

OpenStudy (anonymous):

idk, cuz i got 12

OpenStudy (phi):

this is a bit tricky, notice the theorem says if it is divided by (x - a), the remainder is f(a). you have (x+3) which you should write it as (x - (-3) ) in other words, a is -3

OpenStudy (vera_ewing):

Ok so what did you get for the remainder? And am I right?

OpenStudy (phi):

you have to redo your calculation with -3 (you used +3, which gives 12)

OpenStudy (anonymous):

that makes more sense

OpenStudy (vera_ewing):

Ohhhh ok so it IS 96! right? @phi

OpenStudy (anonymous):

yea it is, i just re done it with -3

OpenStudy (anonymous):

sorry for the confusion @vera_ewing